# The Rate-Distortion-Perception-Classification Tradeoff: Joint Source Coding and Modulation via Inverse-Domain GANs

### v1

@@ IV-C Solving RDPC: Heuristic Algorithm
RDPCO algorithm. We summarize all the above steps in Algorithm 2, which we name RDPCO for RDPC optimization. In our implementation and experiments, we set $p_{0}=p_{1}=1/2$ and, in step 12, employ a constant learning rate of $10^{-4}$ for $2\times 10^{4}$ iterations. Regarding the barrier method in steps 14-17, we initialize $t$ as $t_{0}=0.01$ and update it with $\mu=2$. That algorithm stops either when the number of iterations reaches $m/100$ (determined experimentally) or when the duality gap is below $0.01$. The parameter $\epsilon$ in step 19 is set to $10^{-5}$. To balance the terms in (44), we set $\lambda_{D}=1/\log(D)$, $\lambda_{P}=1/\log(P)$, and $\lambda_{C}=-1/\log(\sqrt{p_{0}\,p_{1}})$.
@@
[…]
Figure 3: Architecture of the encoder $e(\cdot\,;\,\bm{\theta}_{e})$, consisting of fully connected layers of indicated dimensions.
V Experimental Results
We now present our experiments to evaluate the performance of the proposed ID-GAN method (Algorithm 1, for the training of the encoder), and also to show how RDPCO (Algorithm 2) unveils insights about encoder unveils insights about the RDPC problem (7).
V-A ID-GAN-based JSCM framework
Before the experimental setup, we explain how all the functions in the ID-GAN framework in Fig. 1 were implemented.
Network architectures. Fig. 2 shows the architectures of the decoder $d(\cdot\,;\,\bm{\theta}_{d})$ [Fig. 2(a)] and of the critics $f_{1}$ and $f_{2}$ [Fig. 2(b)]. The latter have the same architecture, but they are initialized independently, with different seeds. The decoder in Fig. 2(a) increases the dimensions of the data by first using a fully connected network, whose output is reshaped to a $4\times 4\times 256$ tensor, and then by upsampling along the channel. The upsampling is performed via transposed convolutions with ReLU activations. The architecture of the critics [Fig. 2(b)] is symmetric to that of the decoder. The input image is compressed via a convolutional network, whose output at the last layer is mapped to a probability vector with a sigmoid function. The architecture of the encoder, on the hand, downsamples the input image using only fully connected layers, as shown in Fig. 3. The network architecture of the classifier [cf. Fig. 1, bottom] is similar to the one of the encoder, except that we apply batch normalization after each fully-connected layer and use a leaky ReLU as activation. Also, the last layer is mapped to a 10-dimensional vector (coinciding with the number of classes of MNIST) and a sigmoid is applied to each entry. We train the classifier beforehand and fix it when training the encoder.
Experimental setup. To illustrate the performance of the algorithms, we use the popular MNIST dataset [36], which has $60,000$ training images and $10,000$ test images of size $28\times 28$. In Algorithm 1, we used a learning rate $\alpha=10^{-5}$ and acceleration parameters $\beta_{1}=0.5$ and $\beta_{2}=0.9$ for Adam, a batch size $S=50$, $n_{\text{critic}}=5$ iterations for the inner loop of the critic, and loss hyperparameters $\lambda_{g}=\lambda_{d}=\lambda_{p}=\lambda_{c}=1$. Also, we ran the algorithm for just $5$ epochs. The reason for such a small number is that, as described in Section IV-A, the decoder has already been trained when we run Algorithm 1. And the decoder required also only $8$ epochs to train.
Refer to captionMean-squared error (MSE)Rate
Refer to captionFréchet inception distance (FID)Rate
Refer to captionClassification errorRate
Figure 4: Results for proposed ID-GAN (under different hyperparameters), D-JSCC [2], and AE+GAN training [8]. (a) mean-squared error (MSE), (b) Fréchet inception distance (FID), and (c) classification error as a function of the rate.
[…]
Figure 5: Example images reconstructed by a traditional method (JPEG+LDPC+BPSK), D-JSCC, AE+GAN training, and proposed ID-GAN.
[…]
Figure 6: (a) Rate-distortion curves of RDPCO for varying $P$ and $C$, (b) Corresponding behavior of ID-GAN on MNIST images. The triples on top of each column represent the hyperparameters $(\lambda_{d},\lambda_{p},\lambda_{c})$.
@@ V-A ID-GAN-based JSCM framework
Algorithms. We compare the proposed method for JSCM, i.e., ID-GAN in Algorithm 1, against D-JSCC [2], the parallel autoencoder-GAN (AE+GAN) training method [8], and a traditional approach in which source coding and modulation are designed separately using JPEG and Huffman codes for source coding, $3/4$ LDPC codes for channel coding, and BPSK for modulation [37].
Metrics and comparison. For comparison metrics, we selected the mean squared error (MSE), the Fréchet inception distance (FID) [38], and classification error. FID captures perception quality by measuring the similarity between distributions of real and generated images. The smaller the FID, the closer both distributions are. In summary, the smaller all the metrics, the better the performance of the algorithm. Comparing the performance of a JSCM system against a traditional system, however, is challenging. For example, [2] proposed to use the ratio of bandwidth compression. Yet, in a traditional system, it is not obvious how to accurately determine the ratio between the size of images and the corresponding vectors in IQ-domain. Instead, we will use rate as defined by the objective of (7), which quantifies the amount of information that a symbol can convey through a channel. In particular, the dimension $m$ of the latent variable $\bm{y}$ corresponds to the number of constellations in a traditional system. To compare the different algorithms in a fair way, we thus fix rate, which is independent of the latent/compressed dimension $m$.
Results. Figs. 4-4 show how the above metrics vary with the rate for the proposed ID-GAN with $\lambda_{c}=30$, $\lambda_{c}=100$, and $\lambda_{d}=10$ (when one parameter is set, the remaining are equal to $1$), as well as for D-JSCC [2] and AE+GAN [8]. Fig. 4 indicates that image distortion, measured by the MSE, decreases with transmission rate for all the algorithms. D-JSCC, however, outperforms all other methods, as it was designed to minimize image distortion. Indeed, according to the RDPC tradeoff (Theorem 1), if perception and classification metrics are ignored, an algorithm can achieve a given level of distortion with a smaller transmission rate. The RDPC tradeoff also explains why increasing $\lambda_{d}$ in ID-GAN yields a better MSE [the weight associated to the MSE loss in (15) increases], while increasing $\lambda_{c}$ leads to a worse one. Fig. 4 shows that our algorithm outperforms D-JSCC and AE+GAN in terms of FID for all the values of the tested hyperparameters. This is because we use a GAN to match the distributions of real and generated images. Notice also that in our algorithm FID is essentially invariant to different hyperparameters, since the decoder/generator is fixed in the second stage of training [cf. Fig. 1]. However, when the transmission rate is small, from 0 to 10, FID in ID-GAN decreases drastically from 110 to 80. We attribute this behavior to a relatively large noise in the channel when the rate is small. As the decoder is fixed, a large noise in the output of the encoder $\bm{y}$ translates into images with low-quality in terms of perception. Fig. 4 illustrates again the RDPC tradeoff: when we increase $\lambda_{d}$ in our algorithm, its classification performance decreases. On the other hand, as expected, increasing $\lambda_{c}$ improves its classification performance.
In conclusion, these results indicate that by adopting a two-stage training process whereby the decoder is first trained and then kept fixed, we effectively stabilize the perception quality of the images. Compared to D-JSCC [2], this, however, comes at the expense of a slight increase in MSE.

### v2

@@ IV-B RDPCO: Heuristic algorithm for RDPC optimization
RDPCO algorithm. We summarize all the above steps in Algorithm 1, which we name RDPCO for RDPC Optimization. Steps 5-8 describe the procedure to generate the covariance matrix $\widehat{\bm{\Sigma}}=\bm{D}(\bm{E}\bm{E}^{\top}+\bm{\Sigma})\bm{D}^{\top}$, whose factors are then computed in steps 11-22. The barrier method in steps 14-17 stops whenever the duality gap is below $0.01$ or the number of iterations reaches $m/100$, both parameters determined experimentally. The remaining parameters that we used in our experiments are reported in Section VI-A.
V Solving RDPC with Inverse-Domain GAN
RDPCO attempts to solve the RDPC problem (7) under restrictive assumptions [see Assumption 2]. In this section, leveraging the modeling power of neural networks, in particular generative adversarial networks (GANs) [goodfellow2014generative], we propose an algorithm that works under more general assumptions. In our channel diagram (2), we will thus model the encoder $p_{\bm{Y}|\bm{X}}$ with a neural network $e(\cdot\,;\bm{\theta}_{e})\,:\,\mathbb{R}^{n}\to\mathbb{R}^{m}$ parameterized by $\bm{\theta}_{e}$, and the decoder $p_{\widehat{\bm{X}}|\widehat{\bm{Y}}}$ as neural network $d(\cdot\,;\bm{\theta}_{d})\,:\,\mathbb{R}^{m}\to\mathbb{R}^{n}$ parameterized by $\bm{\theta}_{d}$. These networks will be trained as in ID-GAN [IndomainGAN] which, however, was proposed for a task different from JSCC/JSCM. Specifically, given an (adversarially-trained) image generator, the goal in [IndomainGAN] was to train an encoder to obtain a semantically-meaningful latent code for image editing. We adopt this process of training the generator first, and then the encoder.
V-A Proposed scheme
Fig. 1 shows our framework based on ID-GAN. As in [IndomainGAN], we first train an image generator/decoder $d(\cdot\,;\,\bm{\theta}_{d})$ (Fig. 1, top) adversarially against discriminator $f_{1}$, which learns to distinguish a real signal from a randomly generated one, $d(\bm{Z}\,;\,\bm{\theta}_{d})$, where $\bm{Z}\sim\mathcal{N}(\bm{0}_{m},\,\bm{I}_{m})$ is a vector of i.i.d. standard Gaussians. This is the conventional GAN setup [goodfellow2014generative, WassersteinGAN]. As the discriminator is a particular case of a classifier, outputting just a binary signal, it is also known as a critic. Once the decoder is trained, we fix it and train the encoder $e(\cdot\,;\,\bm{\theta}_{e})$ together with its own critic $f_{2}$, which again learns to distinguish real signals from randomly generated from ones (Fig. 1, bottom). Comparing (2) and Fig. 1 (bottom), we see that $p_{\bm{Y}|\bm{X}}$ is implemented by $e(\cdot\,;\,\bm{\theta_{e}})$, $p_{\widehat{\bm{X}}|\widehat{\bm{Y}}}$ is implemented by $d(\cdot\,;\,\bm{\theta_{d}^{\star}})$, and the Gaussian channel noise $p_{\widehat{\bm{Y}}|\bm{Y}}(\widehat{\bm{y}}|\bm{y})=\mathcal{N}(\bm{y},\bm{\Sigma})$ has (diagonal) covariance matrix $\bm{\Sigma}=\sigma_{t}^{2}\bm{I}_{m}$, where $\sigma_{t}$ is a parameter we learn (or fix) during training. If we normalize the output $\bm{Y}$ of the encoder to have unit power, then the signal-to-noise ratio (SNR) is determined by $\sigma_{t}$ as $\text{SNR}_{t}=-10\log_{10}\,\sigma_{t}^{2}$, and the channel rate is
$R=m\log_{2}\Big{(}1+\frac{1}{\sigma_{t}^{2}}\Big{)}\,.$ (41)
Next, we explain the training processes of the decoder and encoder in more detail.
@@ V-A Proposed scheme
Figure 1: Proposed ID-GAN framework for solving the RDPC problem (7). The decoder is first trained adversarially with critic 1 in the first step (top). The decoder is then fixed and coupled with an encoder, which is in turn trained with critic 2 in order to preserve both reconstruction quality and classification accuracy (bottom). Critics 1 and 2 have the same architecture.
V-B Training the decoder
To train the decoder $d(\cdot\,;\,\bm{\theta}_{d})$ and critic $f_{1}$ as in Fig. 1 (top), we use the Wasserstein GAN (WGAN) [WassersteinGAN] framework. This consists of finding the parameters $\bm{\theta}_{d}$ of the decoder that minimize the Wasserstein-1 (or earth-mover) distance $W_{1}(p_{r},p_{\bm{\theta}_{d}})$ between the distribution $p_{r}$ of real data and the distribution $p_{\bm{\theta}_{d}}$ of data generated by $d(\bm{Z}\,;\,\bm{\theta}_{d})$, with $\bm{Z}\sim\mathcal{N}(\bm{0}_{m},\,\bm{I}_{m})$. We consider the Kantarovich-Rubinstein dual form of the Wasserstein-1 distance:
where the supremum is over the functions $f_{1}\,:\,\mathbb{R}^{n}\to\mathbb{R}$ that are $1$-Lipschitz continuous. While the critic $f_{1}$ is found by maximizing the argument in (42), the parameters of the decoder are found by minimizing the full Wasserstein distance $W_{1}(p_{r},p_{\bm{\theta}_{d}})$. This distance enables overcoming the mode collapse observed in the original GAN framework [goodfellow2014generative], which used instead the Jensen-Shannon divergence. To enforce Lipschitz-continuity of the critic $f_{1}$, [WassersteinGAN] proposed to limit its parameters to a small box around the origin. The work in [WGAN-GP], however, found that this technique leads to instabilities in training (exploding/vanishing gradients) and showed that a gradient penalty solves these problems. We thus adopt the loss suggested in [WGAN-GP] for finding critic $f_{1}$:
where $\lambda_{g}\geq 0$, and $\widetilde{\bm{X}}$ is a point sampled uniformly on the line joining a real data point $\bm{X}\sim p_{r}$ and a point $\bm{Y}\sim p_{\bm{\theta}_{d}}$ generated by the decoder. The third term in (43) eliminates the need to constrain the critic to be Lipschitz-continuous [constraint in (42)]; see [WGAN-GP] for more details. In turn, the parameters $\bm{\theta}_{d}$ of the decoder are found by minimizing (42) which, when $f_{1}$ is fixed, is equivalent to minimizing
$L_{\bm{\theta}_{d}}=-\mathbb{E}_{\bm{X}\sim p_{\bm{\theta_{d}}}}[f_{1}(\bm{X})]=-\mathbb{E}_{\bm{Z}\sim\mathcal{N}(\bm{0}_{m},\,\bm{I}_{m})}\big{[}f_{1}\big{(}d(\bm{Z}\,;\,\bm{\theta}_{d})\big{)}\big{]}.$ (44)
During training, there are two nested loops: the outer loop updates $\bm{\theta_{d}}$; and the inner loop, which runs for $n_{\text{critic}}$ iterations, updates the parameters of the critic such that the supremum in (42) is reasonably well computed. See [WassersteinGAN, WGAN-GP] for details.
Algorithm 2 ID-GAN compression: training of the encoder
1:Training images/labels $\{(\bm{x}^{(t)},\bm{\ell}^{(t)})\}_{t=1}^{T}$, pretrained decoder $d(\cdot)$, pretrained classifier $c(\cdot)$, learning rate $\alpha$, momentum parameters $\beta_{1}$, $\beta_{2}$, batch size $S$, number of iterations of critic $n_{\text{critic}}$, and loss hyperparameters $\lambda_{g}$, $\lambda_{d}$, $\lambda_{p}$, $\lambda_{c}$
2:Set encoder $\bm{\theta}_{e}^{(1)}$ and critic $\bm{\theta}_{f}^{(1)}$ parameters randomly; set channel noise standard deviation $\sigma_{t}^{(1)}>0$ randomly
3:
4:
5:$\mathcal{S}=\text{randperm}(\{1,2,\ldots,T\})$
6:for $j=1,\ldots,\lceil T/S\rceil$ do
7:    Select next $S$ batch indices $\mathcal{S}_{j}$ from $\mathcal{S}$
8:    for $k=1$ to $n_{\text{critic}}$ do
9:         Generate (channel) noise $\bm{Z}\sim\mathcal{N}(\bm{0}_{m},\bm{I}_{m})$
10:         $L_{f_{2}}^{(1)}=\frac{1}{S}\sum_{s\in\mathcal{S}_{j}}f_{2}\Big{(}d\big{(}e(\bm{x}^{(s)}\,;\,\bm{\theta}_{e}^{(j)})+\sigma_{t}^{(j)}\bm{Z}\big{)}\,;\,\bm{\theta}_{f}^{(k)}\Big{)}$
11:         $L_{f_{2}}^{(2)}=\frac{1}{S}\sum_{s\in\mathcal{S}_{j}}f_{2}\Big{(}\bm{x}^{(s)}\,;\,\bm{\theta}_{f}^{(k)}\Big{)}$
12:         for $s\in\mathcal{S}_{j}$ do
13:             Draw $\epsilon\sim\mathcal{U}(0,1)$ randomly
14:             $\widetilde{\bm{x}}^{(s)}=(1-\epsilon)\cdot\bm{x}^{(s)}+\epsilon\cdot d\big{(}e\big{(}\bm{x}^{(s)}\,;\,\bm{\theta}_{e}^{(j)}\big{)}+\sigma_{t}^{(j)}\bm{Z}\big{)}$
15:         end for
16:         $L_{f_{2}}^{(3)}=\frac{1}{S}\sum_{s\in\mathcal{S}_{j}}\Big{(}\big{\|}\nabla_{\widetilde{\bm{x}}}f_{2}(\widetilde{\bm{x}}^{(s)})\big{\|}_{2}-1\Big{)}^{2}$
17:         $L_{f_{2}}=L_{f_{2}}^{(1)}-L_{f_{2}}^{(2)}+\lambda_{g}L_{f_{2}}^{(3)}$
18:         $\bm{\theta}_{f}^{(k+1)}=\text{Adam}\Big{(}\bm{\theta}_{f}^{(k)},\,\lambda_{p}L_{f_{2}},\,\alpha,\,\beta_{1},\,\beta_{2}\Big{)}$
19:    end for
20:    $\bm{\theta}_{f}^{(j)}=\bm{\theta}_{f}^{(n_{\text{critic}})}$
21:    for $s\in\mathcal{S}_{j}$ do
22:         $\widehat{\bm{x}}^{(s)}=d\Big{(}e\big{(}\bm{x}^{(s)}\,;\,\bm{\theta}_{e}^{(j)}\big{)}+\sigma_{t}^{(j)}\bm{Z}\Big{)}$, w/ $\bm{Z}\!\sim\!\mathcal{N}(\bm{0}_{m},\bm{I}_{m})$
23:    end for
24:
$L_{e}=m\log_{2}\Big{(}1+1/{\sigma_{t}^{(j)}}^{2}\Big{)}+\frac{1}{S}\sum_{s\in\mathcal{S}_{j}}\lambda_{d}\Big{\|}\bm{x}^{(s)}-\widehat{\bm{x}}^{(s)}\Big{\|}_{2}^{2}\\[-5.69046pt] +\lambda_{c}\text{CE}\big{(}c(\widehat{\bm{x}}^{(s)}),\bm{\ell}^{(s)}\big{)}-\lambda_{p}f_{2}\Big{(}\widehat{\bm{x}}^{(s)}\,;\,\bm{\theta}_{f}^{(j)}\Big{)}$
25:     $\Big{(}\bm{\theta}_{e}^{(j+1)},\sigma_{t}^{(j+1)}\Big{)}=\text{Adam}\Big{(}\big{(}\bm{\theta}_{e}^{(j)},\sigma_{t}^{(j)}\big{)},\,L_{e},\,\alpha,\,\beta_{1},\,\beta_{2}\Big{)}$
26:end for
V-C Training the encoder
After training the decoder, we fix its parameters to $\bm{\theta}_{d}^{\star}$ and consider the scheme in Fig. 1 (bottom) to train the encoder $e(\cdot\,;\,\bm{\theta}_{e})$. As the decoder, the encoder is also trained adversarially against a critic $f_{2}$ to enhance perception quality, but also takes into account the reconstruction quality and semantic meaning of the reconstruction. The former is captured by an MSE loss between the original and reconstructed images, and the latter by a cross-entropy loss between the image label and the output of a pre-trained classifier $c(\cdot)$ applied to the reconstructed image.
Derivation of the loss for the encoder. To motivate our loss for the encoder, we start from the RDPC problem (7), considering $\Delta(\bm{x},\bm{y})=\|\bm{x}-\bm{y}\|_{2}^{2}$ as the distortion metric, $d(p_{\bm{X}},p_{\widehat{\bm{X}}})=W_{1}(p_{\bm{X}},p_{\widehat{\bm{X}}})$ as the perception metric, and the cross-entropy $\epsilon_{c_{0}}(\bm{X},\widehat{\bm{X}})=\text{CE}(c(\widehat{\bm{X}}),\ell(\bm{X}))$ as the classification loss, where $\ell(\bm{X})$ denotes the class of $\bm{X}$. Then, there exist constants $\lambda_{d}$, $\lambda_{p}$, and $\lambda_{c}$ (related to $D$, $P$, and $C$), such that (7) and
$\underset{p_{\bm{Y}|\bm{X}},\,p_{\widehat{\bm{X}}|\widehat{\bm{Y}}},\,\bm{\Sigma}}{\min}\,\,\,\sum_{i=1}^{m}\log\Big{(}1+\frac{1}{\bm{\Sigma}_{ii}}\Big{)}+\mathbb{E}\Big{[}\lambda_{d}\big{\|}\bm{X}-\widehat{\bm{X}}\big{\|}_{2}^{2}\\ +\lambda_{c}\text{CE}\big{(}c(\widehat{\bm{X}}),\,\ell\big{(}\bm{X}\big{)}\big{)}\Big{]}+\lambda_{p}W_{1}\big{(}p_{\bm{X}},\,p_{\widehat{\bm{X}}}\big{)}$ (45)
have the same solution. Problem (45) is non-parametric, i.e., the functions representing the encoder $p_{\bm{Y}|\bm{X}}$ and the decoder $p_{\widehat{\bm{X}}|\widehat{\bm{Y}}}$ have no structure. As mentioned, we assume they are implemented by neural networks $e(\cdot\,;\bm{\theta_{e}})$ and $d(\cdot\,;\bm{\theta}_{d})$, respectively. The encoder, in particular, normalizes its output signal to unit power. According to the channel model (2), the output of the decoder is then $\widehat{\bm{X}}=d(e(\bm{X})+\bm{N})$, where we omitted dependencies on $\bm{\theta}_{e}$ and $\bm{\theta}_{d}$ for simplicity. Under these assumptions and further assuming that the different channels have equal variance [i.e., (41)], (45) becomes
$\underset{\bm{\theta}_{e},\sigma_{t}}{\text{minimize}}\,\,\,m\log_{2}\Big{(}1+\frac{1}{\sigma_{t}^{2}}\Big{)}+\mathbb{E}\Big{[}\lambda_{d}\big{\|}\bm{X}-d(e(\bm{X})+\sigma_{t}\bm{Z})\big{\|}_{2}^{2}\\ +\lambda_{c}\text{CE}\Big{(}c\big{(}d(e(\bm{X})+\sigma_{t}\bm{Z})\big{)},\,\ell\big{(}\bm{X}\big{)}\Big{)}\Big{]}+\lambda_{p}W_{1}\big{(}p_{\bm{X}},\,p_{\widehat{\bm{X}}}\big{)}\,,$ (46)
where, akin to the reparameterization trick [Kingma14-AutoEncodingVariationalBayes], we replaced $\bm{N}$ by $\sigma_{t}\bm{Z}$, with $\bm{Z}\sim\mathcal{N}(\bm{0}_{m},\bm{I}_{m})$. This makes the dependency of $\bm{N}$ on $\sigma_{t}$ explicit and enables computing derivatives with respect to $\sigma_{t}$. Note that the expectation in (46) is with respect to $\bm{X}$ and $\bm{Z}$. Note also that $\bm{\theta}_{d}$ is not included in the optimization variables of (46), as the decoder has already been trained. Adopting the approximations for $W_{1}$ described in Section V-B, we find the parameters of the encoder and channel noise level by solving
$\underset{\bm{\theta}_{e},\sigma_{t}}{\text{minimize}}\,\,\,m\log_{2}\Big{(}1+\frac{1}{\sigma_{t}^{2}}\Big{)}+\mathbb{E}\Big{[}\lambda_{d}\big{\|}\bm{X}-d(e(\bm{X})+\sigma_{t}\bm{Z})\big{\|}_{2}^{2}\\ +\lambda_{c}\text{CE}\Big{(}c\big{(}d(e(\bm{X})+\sigma_{t}\bm{Z})\big{)},\,\ell\big{(}\bm{X}\big{)}\Big{)}\\ -\lambda_{p}f_{2}(d(e(\bm{X})+\sigma_{t}\bm{Z}))\Big{]}\,.$ (47)
As remarked in Section I-A, we design the encoder and channel noise level $\sigma_{t}$ jointly. In practice, one would instead estimate the channel noise level and design the power of the encoder. The advantage of doing as we do is that the first term of (47) is convex in $\sigma_{t}$. The parameters $\bm{\theta}_{f_{2}}$ of critic $f_{2}$, in turn, are computed like in (43), by minimizing the loss
$L_{f_{2}}=\lambda_{p}\bigg{\{}\mathbb{E}\Big{[}f_{2}\big{(}d(e(\bm{X})+\sigma_{t}\bm{Z})\big{)}-f_{2}(\bm{X})\Big{]}\\ +\lambda_{g}\,\mathbb{E}\Big{[}\big{(}\big{\|}\nabla_{\widetilde{\bm{x}}}f_{2}\big{(}\widetilde{\bm{X}}\big{)}\big{\|}_{2}-1\big{)}^{2}\Big{]}\bigg{\}}\,,$ (48)
where $\widetilde{\bm{X}}=(1-\epsilon)\bm{X}+\epsilon\,d(e(\bm{X})+\sigma_{t}\bm{Z})$ and $\epsilon\sim\mathcal{U}(0,1)$ is uniformly distributed in $[0,1]$. The first expectation is with respect to $\bm{X}$ and $\bm{Z}$, and the second with respect to $\bm{X}$ and $\epsilon$.
Training algorithm. The complete training procedure of the encoder is shown in Algorithm 2. Its inputs include training images $\bm{x}^{(t)}$ and corresponding labels $\bm{\ell}^{(t)}$, and a pretrained decoder $d(\cdot)$ and classifier $c(\cdot)$. After initializing the parameters of the encoder, associated critic, and channel noise level, in each epoch we randomly permute the indices of the training data (step 5) and visit all the training data in batches of size $S$. This takes $\lceil T/S\rceil$ iterations, where $T$ is the number of data points. The loop in steps 8-19 performs $n_{\text{critic}}$ iterations of Adam to minimize the critic loss in (48) and thus to update the critic $f_{2}$ parameters $\bm{\theta}_{f}$ (where we omit the index $2$ for simplicity). This corresponds to computing the supremum in the Wasserstein distance (42) between the real data $\bm{X}$ and the reconstructed one $\widehat{\bm{X}}=d\big{(}e\big{(}\bm{X}\big{)}+\bm{N}\big{)}$. The terms in (43) are computed separately, with the last term requiring the creation of the intermediate variables $\widetilde{\bm{x}}^{(s)}$ in steps 12-15. As usual, expected values were replaced by sample averages over the batch. After having updated the parameters of critic $f_{2}$, we perform one iteration of Adam to minimize the encoder loss in (47) and thus to update the encoder parameters $\bm{\theta}_{e}$ and channel noise level $\sigma_{t}$. This requires passing each image in the batch through the encoder and decoder to create $\widehat{\bm{x}}^{(t)}$, as in steps 21-23. In step 25, the parameters of the encoder and the channel noise level are updated simultaneously. In our experiments, reported in Section VI, we run two versions of Algorithm 2: one exactly as described, where the noise level $\sigma_{t}$, and thus $\text{SNR}_{t}=-10\log_{10}\sigma_{t}^{2}$, is learned during training; and another where $\text{SNR}_{t}$ is fixed to a predefined value.
@@ V-C Training the encoder
VI Experimental Results
We now present our experiments to evaluate the performance of the proposed algorithms, RDPCO (Algorithm 1) and ID-GAN (Algorithm 2). We start with RDPCO and then consider ID-GAN.
VI-A RDPCO algorithm
Recall that RDPCO (Algorithm 1) solves the approximation (32) of (7). Before explaining the experiments, we describe how we set the parameters of the algorithm.
Experimental setup. In Algorithm 1, we generate $\bm{c_{n}}\in\mathbb{R}^{n}$ randomly with i.i.d. Gaussian entries with zero mean and variance $4$. The classes are always equiprobable, i.e., $p_{0}=p_{1}=1/2$. In the gradient descent method in step 12, we employ a constant learning rate of $10^{-4}$ for $20$k iterations. In the barrier method in steps 14-17, we initialize $t$ as $t_{0}=0.01$ and update it with a factor of $\mu=2$. The parameter $\epsilon$ in step 19 is set to $10^{-5}$. To balance the terms in (39), we set $\lambda_{D}=1/\log(D)$, $\lambda_{P}=1/\log(P)$, and $\lambda_{C}=-1/\log(\sqrt{p_{0}\,p_{1}})$. During the experiments, we vary $D$, $P$, $C$, and $R$ [which depends on the latent dimension $m$ and noise level $\sigma_{t}$; cf. (41)]. To evaluate the performance of the algorithm, we visualize how two metrics vary, e.g., rate and distortion, while the remaining parameters are fixed.
Metrics as a function of $\bm{D}$. Here, we fix the input dimension to $n=5$ and the latent one to $m=2$. Fig. 2 shows how distortion, perception, and classification error metrics vary with $D$ in (32). These metrics are, respectively, the right-hand side of (22), of (24), and of (26). In Figs. 22, we see that when $C$ (resp. $P$) is fixed, increasing $P$ (resp. $C$) increases either the distortion or perception metrics. This behavior is as expected according to Theorem 1. In Fig. 2, we observe that for a fixed $C$, modifying $P$ produces no significant effect on the classification error, indicating that the classification constraint becomes active before the perception one.
Fig. 3 is similar to Fig. 2, but $P$ and $C$ are fixed to $4.1$ and $0.1$, respectively, while the latent dimension $m$ varies. When $m=1$, all metrics are invariant to $D$. This is because, as seen in Fig. 3, the perception constraint is active (it equals its limit of $4.1$), dominating the two other constraints. For $m=2,3$, their classification error behaves similarly, but $m=3$ achieves better perception and worse distortion.